OpenStudy (anonymous):

Help!!! (exponential growth problem) The earth has 3.2x10^9 acres of arable land available. The world population of 1950 required 10^9 acres to sustain it, and the population of 1980 required 2x10^9 acres. If the required acreage grows at a constant percentage rate, in what year will the population reach the maximum sustainable size?

OpenStudy (anonymous):

It takes 30 years for the things to double.

OpenStudy (anonymous):

Since we are given a doubling of the area... we'll let 2 be our base. \[ A(t) = C_02^{kt} \]It takes 30 years to double, so when \(t=30\) we want \(kt=1\). \[ 1 = k(30) \implies k= \frac{1}{30} \]

OpenStudy (anonymous):

Lastly we must solve for \(C_0\), given the fact that \(A(1950)=10^{9}\): \[ \large A(1950) = C_0 2^{1950/30} = 10^9 \]

OpenStudy (anonymous):

so once we have this how do we determine when the population will reach the max sustainable size

OpenStudy (anonymous):

You solve for \(t\) when \(A(t) = 3.2 \times 10^{9}\)

OpenStudy (anonymous):

A(t)=3.2×10^9=C02^1950/30 like this???

OpenStudy (anonymous):

You still need to solve for \(C_0\).

OpenStudy (anonymous):

and do i do that by dividing 10^9 by Co^1950/30? cause if I do that I get a really really small number

OpenStudy (anonymous):

......

OpenStudy (anonymous):

You do that by solving for it in the equation I already gave.

OpenStudy (anonymous):

\[ \large A(1950) = C_0 \cdot 2^{1950/30} = 10^9 \]

OpenStudy (anonymous):

I am not sure how to do that. This is question I have seen. I am sorry

OpenStudy (anonymous):

It's called division...

OpenStudy (anonymous):

I understand that but what do I divide?